# CodeUltraFeedback: An LLM-as-a-Judge Dataset for Aligning Large Language Models to Coding Preferences

### v2

@@ 2. CodeUltraFeedback
2.1. Coding Preferences and Principles.
@@
2.3. LLMs Selection.
In order to generate highly diverse solutions to the instructions, we select a pool of 14 LLMs spanning eight model families. These include GPT-4-Turbo and GPT-3.5-Turbo as closed-source LLMs. We include Llama-2-13B-Chat and Llama-2-70B-Chat (Touvron et al., 2023), specialized for chat use cases. We choose CodeLlama-7/13/34B-Instruct models (Roziere et al., 2023), instruction-tuned for code generation. We also include WizardLM-15/33B models (Xu et al., 2023) and Mistral-7B-Instruct (Jiang et al., 2023), which are instruction-tuned LLMs that can follow complex instructions. Finally, we include WizardCoder-15/33B (Luo et al., 2023) and DeepSeek-Coder-6.7/33B-Instruct (Guo et al., 2024), recent instruction-tuned LLMs that demonstrate state-of-the-art performance on code generation benchmarks such as HumanEval (Chen and others., 2021) and HumanEval+ (Liu et al., 2023b). From this pool, we randomly select four LLMs for each instruction to generate responses, allowing for the representation of a variety of coding approaches and styles.
2.4. Principle-Driven Code Generation.
@@ 3.1. LLMs Scores Exploration
At first glance, the scores might suggest marginal discrepancies given the 1–5 scoring range. However, these variations are substantial and carry statistical significance. To quantify this significance, we performed pairwise Welch’s $t$-tests, comparing the score distribution for each LLM against GPT-3.5-Turbo across each coding preference. The results reveal highly significant statistical differences, with $p<0.001$, for most comparisons and validate the discernible performance gaps between all models and GPT-3.5-Turbo. Moreover, the results also demonstrate a notably superior performance of GPT-4-Turbo relative to GPT-3.5-Turbo.
@@ 6. Discussion
In this paper, we have demonstrated how LLM-as-a-Judge enables the creation of new datasets and benchmarks like CodeUltraFeedback and CodeUltraFeedback-Bench to facilitate the tuning of LLMs for better alignment with coding preferences and assess their performance. This section offers further insights from our experiments, potential improvements for LLM alignment, and the significance of exploring varied LLM-as-a-Judge configurations. Additionally, we outline some limitations of CodeUltraFeedback and CodeUltraFeedback-Bench.
@@ 6.2. Limitations
Our methodology presupposes the existence of a highly capable model, such as GPT-3.5 or GPT-4, as a proxy to humans to accurately evaluate other LLMs. The reliance on these models is predicated upon previous findings, which indicate that GPT-3.5 and GPT-4 align closely with human judgments (Zheng et al., 2023a; Li et al., 2023b), achieving agreement rates on par with those between humans themselves on MT-Bench (Zheng et al., 2023a). Additionally, recent work (Zhuo, 2024) demonstrates that ICE-Score, a novel metric for code generation based on LLM-as-a-Judge is superior to prior metrics and has a higher correlation with human judgements. Such evidence supports the utilization of LLMs as evaluators. Although our work does not include a comparison between GPT-3.5/4 and human judgements, we hypothesize their correlation on CodeUltraFeedback-Bench to be high and leave the comparison for future work.
@@
Lastly, our grading procedure in CodeUltraFeedback-Bench necessitates reference responses. This usage of reference responses gives the judge LLMs a consistent point of comparison, enabling fair and consistent judgements across LLMs outputs. Future work might explore ways to reduce this dependency.

### v3

@@ 2. CodeUltraFeedback
2.1. Coding Preferences and Principles
@@ 2.2. Initial Dataset
2.3. LLMs Selection
In order to generate highly diverse solutions to the instructions, we select a pool of 14 LLMs spanning eight model families. These include GPT-4-Turbo and GPT-3.5-Turbo as closed-source LLMs. We include Llama-2-13B-Chat and Llama-2-70B-Chat (Touvron et al., 2023), specialized for chat use cases. We choose CodeLlama-7/13/34B-Instruct models (Roziere et al., 2023), instruction-tuned for code generation. We also include WizardLM-15/33B models (Xu et al., 2023) and Mistral-7B-Instruct (Jiang et al., 2023), which are instruction-tuned LLMs that can follow complex instructions. Finally, we include WizardCoder-15/33B (Luo et al., 2023) and DeepSeek-Coder-6.7/33B-Instruct (Guo et al., 2024), recent instruction-tuned LLMs that demonstrate state-of-the-art performance on code generation benchmarks such as HumanEval (Chen and others., 2021) and HumanEval+ (Liu et al., 2023b). From this pool, we randomly select four LLMs for each instruction to generate responses, allowing for the representation of a variety of coding approaches and styles. By including models that encompass diverse pre-training and instruction-tuning phases, our dataset captures a wide spectrum of response generation behaviors. The differences between LLMs extend to pre-training objectives, with some models specializing in general natural language tasks (e.g., Llama-2-13B-Chat) and others explicitly optimized for code-related tasks (e.g., WizardCoder and DeepSeek-Coder). This diversity ensures that the dataset reflects not only distinct model capabilities but also various stylistic and structural approaches to solving programming problems. In turn, this strategy prevents subsequent fine-tuning processes that use the CodeUltraFeedback preference dataset to overfit a specific style of response.
2.4. Principle-Driven Code Generation
@@ 3.1. LLMs Scores Exploration
At first glance, the scores might suggest marginal discrepancies given the 1–5 scoring range. However, these variations are substantial and carry statistical significance. To quantify this significance, we performed pairwise Welch’s $t$-tests, comparing the score distribution for each LLM against GPT-3.5-Turbo across each coding preference. This test was chosen because it accounts for differences in sample sizes and variances between groups. For each comparison, the null hypothesis assumes that the means of the two groups are equal. We calculated p-values to assess the statistical significance of observed differences, using a significance threshold of $p<0.05$. The results reveal highly significant statistical differences, with $p<0.001$, for most comparisons and validate the discernible performance gaps between all models and GPT-3.5-Turbo. Moreover, the results also demonstrate a notably superior performance of GPT-4-Turbo relative to GPT-3.5-Turbo.
@@ 5. Results
Figure 6. Examples of solutions generated by CodeLlama-7B-Instruct+SFT+DPO for two coding preferences: code readability and code explanation.
@@
To further illustrate the alignment of CodeLlama-7B-Instruct+SFT+DPO, we highlight two examples of solutions in Fig. 6 for the code readability and explanation preferences. For the code readability problem, the model’s response demonstrates clear organization, use of comments, and adherence to the task requirements, earning high ratings (GPT-4: 10/10, Human-1: 8/10, Human-2: 10/10). Similarly, in the code explanation problem, the model provides a well-structured, detailed breakdown of the algorithm, including logical flow and computational complexity, achieving a perfect score from all evaluators (GPT-4: 10/10, Human-1: 10/10, Human-2: 10/10). These examples highlight the high-quality outputs of CodeLlama-7B-Instruct+SFT+DPO and its ability to generate solutions tailored to different coding preferences.
@@
LLM Judgements Correlate with Human Evaluators. We investigate the judging capabilities of GPT-3.5 and GPT-4 by evaluating their alignment with human evaluators. We randomly select 20 samples per coding preference, manually annotated by two authors of the paper following the same judging setup as LLM-as-a-Judge. Fig. 8(a) reports Cohen’s kappa coefficients ($\kappa$) for each LLM-human combination. Given that ratings are on a 1-10 scale, we use a tolerance threshold of $\tau=2$, considering ratings like 7 and 9 to be in agreement. We also compare the score distributions of humans and LLMs in Fig. 8(b).
We observe that GPT-3.5 generally shows stronger agreement with human evaluators (H-1 and H-2) across coding preferences compared to GPT-4. For example, GPT-3.5 demonstrates almost perfect agreement with both humans in code readability ($\kappa=0.83$ and $\kappa=0.91$), whereas GPT-4 demonstrates only moderate agreement in this preference ($\kappa=0.48$ and $\kappa=0.43$). Interestingly, for most coding preferences—except for instruction following—GPT-3.5’s judgments show stronger alignment with individual human judges than the agreement observed between the human judges themselves. This observation suggests that GPT-3.5’s evaluations may serve as a reliable proxy for human judgment in these coding preferences.
The score distributions in Fig. 8(b) reveal that humans tend to give higher ratings compared to GPT-3.5 and GPT-4. Notably, GPT-3.5 shows better consistency with human ratings’, while GPT-4’s ratings are more evenly distributed, which contributes to its lower kappa values. Overall, our findings highlight that GPT-3.5 aligns more closely with human evaluators, though both LLMs would benefit from further calibration, particularly in the instruction-following preference.
Preference G-3.5 G-4 H-1 H-2
Instruction-Following G-3.5 - 0.41 0.29 0.32
G-4 - - 0.39 0.43
H-1 - - - 0.57
H-2 - - - -
Explanation G-3.5 - 0.60 0.78 0.67
G-4 - - 0.49 0.62
H-1 - - - 0.64
H-2 - - - -
Efficiency G-3.5 - 0.73 0.50 0.58
G-4 - - 0.49 0.56
H-1 - - - 0.52
H-2 - - - -
Readability G-3.5 - 0.53 0.83 0.91
G-4 - - 0.48 0.43
H-1 - - - 0.64
H-2 - - - -
Style G-3.5 - 0.48 0.77 0.67
G-4 - - 0.48 0.37
H-1 - - - 0.60
H-2 - - - -
(a) Agreement between humans and LLMs judgements in terms of Cohen’s kappa coefficient with a tolerance threshold of $\tau=2$.
(b) Score distribution of humans (H-1 and H-2) and GPT-3.5/4.
Figure 8. Analysis of the judging agreements between humans and LLMs (a) and their score distributions (b) for solutions generated by CodeLlama-7B-Instruct+SFT+DPO.
@@
Figure 9. GPT-3.5 and GPT-4 ratings for HumanEval+ solutions generated by CodeLlama-7B-Instruct+SFT+DPO.
GPT-4 Prefers Test-Passing Solutions. To further analyze the judging behavior of GPT-3.5 and GPT-4, we examine their ratings for solutions generated by CodeLlama-7B-Instruct+SFT+DPO on the HumanEval+ benchmark. Fig. 9 depicts the distribution of ratings assigned by GPT-3.5 (left) and GPT-4 (right), distinguishing between solutions that pass and fail the corresponding test cases. To compute the ratings, we followed an identical LLM-as-a-Judge setup than for CodeUltraFeedback by designing a specific prompt (see Table 17), which focuses on judging functional correctness and the usefulness of the solution.
The results indicate a significant behavioral difference between the GPT-3.5 and GPT-4. GPT-4 shows a clear preference for solutions that pass test cases, assigning them consistently higher ratings. In contrast, GPT-3.5 demonstrates little distinction between test-passing and test-failing solutions, as reflected in the more uniform rating distribution across both categories. These findings underscore GPT-4’s strength in aligning ratings with objective evaluation criteria, such as passing test cases, and highlight the need for further refinement of GPT-3.5 in this regard.
@@ 6. Discussion
In this paper, we have demonstrated how LLM-as-a-Judge enables the creation of new datasets and benchmarks like CodeUltraFeedback and CodeUltraFeedback-Bench to facilitate the tuning of LLMs for better alignment with coding preferences and assess their performance. This section offers further insights from our experiments, potential improvements for LLM alignment and evaluation, and the significance of exploring various LLM-as-a-Judge configurations. Additionally, we outline some limitations of CodeUltraFeedback and CodeUltraFeedback-Bench.
6.1. Limitations of Existing Evaluation Metrics
Evaluation metrics are central to assessing the performance of LLMs for code generation. Automated metrics and static analysis tools provide scalable and objective evaluations. However, while these techniques effectively capture specific aspects of the LLM-generated code, e.g., whether it passes unit tests or matches a ground truth solution, they fail to evaluate the latter against user-centric preferences such as code readability or whether the response strictly follows the user’s instructions.
One of the main limitations of automated metrics is that they cannot assess language subtleties as humans do. For instance, metrics like Pass@$k$ are useful for evaluating functional correctness but cannot discriminate between an overly verbose solution or one lacking meaningful comments and explanations. On the other hand, static analysis tools focus on enforcing rules, which does not encompass any form of language understanding that would allow determining whether a solution includes useful explanations for a given generated code. Altogether, our approach featuring LLM-as-a-Judge offers a solution to these limitations by focusing on LLM alignment with user coding preferences. For example, when evaluating a solution to a coding task, the judge LLM can discern whether comments add meaningful context, whether variable names enhance code readability, or whether an explanation is structured logically. This process is parallel to how humans would judge a solution.
The interplay between LLM-as-a-Judge and automated metrics offers significant opportunities for advancing evaluation methodologies. In particular, the combination of automated evaluation metrics and judging LLMs could result in more holistic evaluations of LLM-generated code and enhance the way we benchmark code LLMs. For instance, future metrics might integrate functional correctness scores with readability and explanation assessments, creating a weighted evaluation system that reflects the multifaceted nature of high-quality code.
@@ 6.3. Limitations
Our methodology presupposes the existence of a highly capable model, such as GPT-3.5 or GPT-4, as a proxy to humans to accurately evaluate other LLMs. The reliance on these models is predicated upon previous findings, which indicate that GPT-3.5 and GPT-4 align closely with human judgments (Zheng et al., 2023a; Li et al., 2023b), achieving agreement rates on par with those between humans themselves on MT-Bench (Zheng et al., 2023a). Additionally, recent work (Zhuo, 2024) demonstrates that ICE-Score, a novel metric for code generation based on LLM-as-a-Judge is superior to prior metrics and has a higher correlation with human judgements. Additionally, we performed a correlation analysis between GPT-3.5, GPT-4, and human evaluators using 100 samples from CodeUltraFeedback-Bench. The results reveal that GPT-3.5 shows a strong correlation with human judgments across coding preferences, while exhibiting a similar score distribution. Such evidence supports the utilization of LLMs as evaluators.
@@
Lastly, our grading procedure in CodeUltraFeedback-Bench necessitates reference responses. This usage of reference responses gives the judge LLMs a consistent point of comparison, enabling fair and consistent judgements across LLMs outputs. Future work might explore ways to alleviate this dependency.
@@ Appendix A Assessment Templates
Table 9. Code readability assessment template.
Scoring: Rate outputs on a scale of 1 to 5:
@@
Table 17. Evaluation template for HumanEval samples.
# Functional Correctness Assessment
Evaluation Criteria:
  - Functional Correctness: Does the code correctly solve the problem as described in the instruction, passing all test cases?
  - Practical Usefulness: Is the response executable and aligned with the intended task requirements?
  - Clarity: Does the code offer a clear and concise solution without unnecessary complexity?
## Format:
### Instruction
  [Instruction]
### Tests:
  [List of unit tests]
### Assistant’s Response:
  [Response]
### Output
Rationale: [Rationale for the rating in short sentences]
Rating: [Rating for Assistant’s Response]
## Examples of output:
Rationale: The code passes all test cases and solves the problem effectively with a concise implementation.
Rating: 10/10 Rationale: The code fails on key test cases and includes unnecessary complexity, making it difficult to understand.
Rating: 3/10 ## Annotation
### Instruction
  {instruction}
### Reference Response:
  {reference}
### Assistant’s Response:
  {response}
### Output